If 9-digit number 2018A902B is divisible by 144, find the value of A+B
Answer by mananth(16946)

If 9-digit number 2018A902B is divisible by 144, find the value of A+B
16*9 =144
the number should be divisible by 16 and 9
Divisibility test for 16
The number has 9 digits
the last 4 digits of the number should be divisible by 16
902B is divisible by 16 if B = 4
if number is divisible by 9 then sum of digits must be 9
plug 4 in B
2018A9024
sum should be 9
The above number sum of digits is 26+A ( 9*3=27)
So A =1
A=1 B=4
A+B = 5
The number is
201819024
201819024 /144 =1401521
It works